If possible, simplify the ratinoal expression.
y^2 - 81 / 2y - 18

Answer by fractalier(6550)

Let us factor top and bottom and see if anything cancels...so from
y^2 - 81 / 2y - 18 =
(y + 9)(y - 9) / 2(y - 9) =
(y + 9) / 2
The y - 9's cancel...